1.113.4

Properly manage the NFS, smb, and nmb daemons

Weight 4

Linux Professional Institute Certification — 102

                Angus Lees gus@inodes.org

Geoffrey Robertson ge@ffrey.com

               Nick Urbanik nicku@nicku.org

2005 July

Description of Objective

Candidate should know how to mount remote filesystems using NFS, configure NFS for exporting local filesystems, start, stop, and restart the NFS server. Install and configure Samba using the included GUI tools or direct edit of the /etc/smb.conf file (Note: this deliberately excludes advanced NT domain issues but includes simple sharing of home directories and printers, as well as correctly setting the nmbd as a WINS client).

Key files, terms, and utilities include:

/etc/exports
/etc/fstab
/etc/smb.conf
mount
umount

\underline{2}

Properly manage the NFS, smb and nmbs daemons

Candidate should know how to mount remote filesystems using NFS, configure NFS for exporting local filesystems, start, stop, and restart the NFS server. Install and configure Samba using the included GUI tools or direct edit of the /etc/smb.conf file (Note: this deliberately excludes advanced NT domain issues but includes simple sharing of home directories and printers, as well as correctly setting the nmbd as a WINS client).

Properly manage the NFS, smb, and nmb daemons

/etc/exports
/etc/fstab
/etc/smb.conf
mount
umount

Properly manage the NFS, smb, and nmb daemons

   NFS:  http://nfs.sourceforge.net/
   Samba:  http://au1.samba.org/samba/docs/

NFS

• “Network File System”

• Extremely widely used protocol for sharing files between Unix machines

• Built over RPC

• Almost always uses UDP

• Unlike alternatives, the NFS protocol is stateless—the server can reboot and the client carries on

• NFSv2 is very simple; v3 is an evolution of v2; v4 is complex (and stateful)

NFS

mount -o options server:/path /mnt_point
# typical /etc/fstab line - see nfs(5)
server:/exports/home /home nfs rw,intr,rsize=8192,wsize=8192

Some options:

rsize/wsize Read/write size

intr Allow signals to interrupt timed-out operations

soft Don’t retry indefinitely NFS

/etc/exports controls what gets exported to where

# /etc/exports.  See exports(5)
/home   10.0.128.0/24(rw,no_root_squash)
/cdrom  *(ro,all_squash)

Information is actually communicated to mountd and kernel NFS server via exportfs(8) Note: No whitespace is allowed between client and client options (showmount(8) will query this information from a server)

Samba

• Implementation of Windows networking protocols (SMB/CIFS, NMB) for Unix

• Samba version 3 (just released) gives PDC/BDC functionality, NT4 domain member support, much improved printing, etc

• Full suite of tools, including smbd and nmbd daemons, but also client tools: smbclient, rpcclient, net, smbstatus, nmblookup, smbtar, smbspool, etc

Samba

   Configured through /etc/samba/smb.conf

win.ini-esque sections describing shares

[public]
comment = Public files
path = /home/ftp/pub
read only = yes
guest ok = yes

Various %X string substitutions available (see smb.conf(5))

Samba

[global] Global configuration

[homes] Triggers creation of automatic user home directory shares

[printers] Triggers creation of printer shares for each system printer

Samba “Samba Web Administration Tool”

   Usual install is http://localhost:901/

Easy to use, links to relevant documentation, etc.

Rewrites smb.conf, loses comments, reorders and destroys hand crafted include, etc directives.

License Of This Document

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005, 2003 Angus Lees <gus@inodes.org>, Geoffrey Robertson <ge@ffrey.com> and Nick Urbanik <nicku@nicku.org>. Permission is granted to make and distribute verbatim copies or modified versions of this document provided that this copyright notice and this permission notice are preserved on all copies under the terms of the GNU General Public License as published by the Free Software Foundation—either version 2 of the License or (at your option) any later version.